OpenStudy (kenhan20):

(−∞,−3] ∪ (4,9] says please enter following interval please help

OpenStudy (danjs):

not sure what you're asking.

zepdrix (zepdrix):

Are you trying to draw the intervals?

OpenStudy (mww):

(a,b ) is the same as a < x < b [a,b] is the same as \[a \le x \le b\] (a,b] is the same as \[a < x \le b\] [a,b) is the same as \[a \le x < b\] So ( or ) does not include the pt, while [ or ] does The U union symbol means 'OR', so you include all of both intervals shown. On the number line